# arXiv:2305.16494v1[cs.CV] 25 May 2023

### v1

@@ Appendix D More Results
Here we present more results randomly sampled from the ImageNet validation set to show that Diff-PGD can steadily generate adv-samples with high stealthiness. In Figure 8, we show some complementary results of $\epsilon=16,\eta=2,n=10$ in main part of the paper. Also in Figure 9, we show results when the perturbation bound is larger: $\epsilon=32,\eta=4,n=10$, where we use DDIM10 with $K_{s}=2$, from the figure we can see that Diff-PGD can still generate realistic adv-samples.
Also, we demonstrate that Diff-PGD remains effective in generating adversarial samples with a high success rate even when the attack bound is smaller. Figure 10 illustrates this point, where we set $\epsilon=8$ and evaluate two Diff-PGD strategies: DDIM50 and DDIM100, all with $T_{s}=2$. Both strategies exhibit a notable success rate and can produce realistic adversarial samples.
Figure 8: More Results of Diff-PGD with $\mathbf{\epsilon=16}$: five columns of each image block follows $x$, $x_{PGD}$, $\delta_{PGD}$, $x^{n}_{0}$ and $\delta^{n}_{0}$ as is defined in the main paper. We can see that Diff-PGD can steadily generate adv-samples with higher stealthiness. Zoom in on a computer screen for better visualization.
Figure 9: More Results of Diff-PGD with $\mathbf{\epsilon=32}$: five columns of each image block represent $x$, $x_{PGD}$, $\delta_{PGD}$, $x^{n}_{0}$ and $\delta^{n}_{0}$ respectively. We can see that Diff-PGD can steadily generate adv-samples with higher stealthiness. Zoom in on a computer screen for better visualization.
Figure 10: Effective of Diff-PGD with Smaller Bound: we show that when the attack $\ell_{\infty}$ bound is smaller: $\epsilon=8$, Diff-PGD is still effective to generate adv-samples with high success rate. (a) we use DDIM100 and $K_{s}=2$, (b) we use we use DDIM50 and $K_{s}=2$.
@@
Diff-rPGD can be adopted in customized settings where only masked regions can be attacked. Here, we show more results of regional attacks in Figure 16, where $\epsilon=16,\eta=2,n=10$. We compare Diff-rPGD with rPGD, and the results show that our Diff-rPGD can deal with the regional adversarial attack better: the masked region can be better merged into the background.
@@
In the main paper, we show that Diff-PGD can help generate style-based adv-samples effectively, in Figure 11 we further present that the second stage in our two-stage strategy is necessary and can highly preserve the realism of the output images.
@@
We present the result of physical world attacks using adversarial patches generated using Diff-PGD in the main paper. In order to show that the adversarial patches are robust to camera views, we show more images taken from different camera views in Figure 12. For the two cases: computer mouse (untargeted) and back bag (targeted to Yorkshire Terrier), we randomly sample ten other camera views. The results show that the adversarial patches are robust both for targeted settings and untargetted settings, for targteted settings, our adversarial patch can fool the network to predict back bag as terriers, and for untargeted settings, the adversarial patch misleads the network to predict computer mouse as artichoke.
@@
We presented the results when $\epsilon=16$ and found that: though designed for improved stealthiness and controllability, it is surprising that Diff-PGD transfers better than the original PGD to five different networks, here we show more results on different $\ell_{\infty}$ bound. We take the other two settings: $\epsilon=8$ and $\epsilon=32$ to better investigate our methods, for $\epsilon=8$ we still use $\eta=2$ while for $\epsilon=32$, we use $\eta=4$ for better attacks, both of them use $n=10$. For Diff-PGD, we use DDIM50 with $K_{s}=2$ for $\epsilon=8$ case and DDIM30 with $K_{s}=2$ for $\epsilon=32$ case.
@@
Figure 13: More Results on Transferability Test: in the main paper we show results of transferability of $\epsilon=16$ settings, here we test transferability of Diff-PGD under different $\ell_{\infty}$ bounds, (Left): $\epsilon=32,\eta=4,n=10$, for Diff-PGD we use DDIM50 with $K_{s}=2$, (Right): $\epsilon=8,\eta=2,n=10$ and for Diff-PGD we use DDIM30 with $K_{s}=2$. From the plots we can see that Diff-PGD has better transferability power than PGD.
Appendix E Human Evaluation
To better evaluate the stealthiness of Diff-PGD compared with the original PGD attack, we conduct a survey among humans with the assistance of Google Form. The user interface of the survey is shown in Figure 17, where the participants are allowed to choose multiple candidate images that they think are realistic. We randomly choose ten images from the ImageNet validation set and the choices for each question include one clean image, one adv-sample generated by PGD, and one adv-sample generated by Diff-PGD. For Diff-PGD we use DDIM50 with $K_{s}=3$, and for both Diff-PGD and PGD we use $\epsilon=16,\eta=2,n=10$.
@@ Appendix E Human Evaluation
Appendix F Potential Social Impacts
@@
Figure 15: More Results of Diff-PGD with $\mathbf{\epsilon=16}$: five columns of each image block follows $x$, $x_{PGD}$, $\delta_{PGD}$, $x^{n}_{0}$ and $\delta^{n}_{0}$ as is defined in the main paper.

### v2

@@ Appendix G More Complementary Results
Here we present more results randomly sampled from the ImageNet validation set to show that Diff-PGD can steadily generate adv-samples with high stealthiness. In Figure 12, we show some complementary results of $\epsilon=16/255,\eta=2/255,n=10$ in main part of the paper. Also in Figure 13, we show results when the perturbation bound is larger: $\epsilon=32/255,\eta=4/255,n=10$, where we use DDIM10 with $K_{s}=2$; from the figure we can see that Diff-PGD can still generate realistic adv-samples.
Also, we demonstrate that Diff-PGD remains effective in generating adversarial samples with a high success rate even when the attack bound is smaller. Figure 14 illustrates this point, where we set $\epsilon=8/255$ and evaluate two Diff-PGD strategies: DDIM50 and DDIM100, all with $T_{s}=2$. Both strategies exhibit a notable success rate and can produce realistic adversarial samples.
Figure 12: More Results of Diff-PGD with $\mathbf{\epsilon=16/255}$: five columns of each image block follows $x$, $x_{PGD}$, $\delta_{PGD}$, $x^{n}_{0}$ and $\delta^{n}_{0}$ as is defined in the main paper. We can see that Diff-PGD can steadily generate adv-samples with higher stealthiness. Zoom in on a computer screen for better visualization.
Figure 13: More Results of Diff-PGD with $\mathbf{\epsilon=32/255}$: five columns of each image block represent $x$, $x_{PGD}$, $\delta_{PGD}$, $x^{n}_{0}$ and $\delta^{n}_{0}$ respectively. We can see that Diff-PGD can steadily generate adv-samples with higher stealthiness. Zoom in on a computer screen for better visualization.
Figure 14: Effective of Diff-PGD with Smaller Bound: we show that when the attack $\ell_{\infty}$ bound is smaller: $\epsilon=8/255$, Diff-PGD is still effective to generate adv-samples with high success rate. (a) we use DDIM100 and $K_{s}=2$, (b) we use we use DDIM50 and $K_{s}=2$.
@@
Diff-rPGD can be adopted in customized settings where only masked regions can be attacked. Here, we show more results of regional attacks in Figure 20, where $\epsilon=16/255,\eta=2/255,n=10$. We compare Diff-rPGD with rPGD, and the results show that our Diff-rPGD can deal with the regional adversarial attack better: the masked region can be better merged into the background.
@@
In the main paper, we show that Diff-PGD can help generate style-based adv-samples effectively. In Figure 15, we further present that the second stage in our two-stage strategy is necessary and can highly preserve the realism of the output images.
@@
We present the result of physical world attacks using adversarial patches generated using Diff-PGD in the main paper. In order to show that the adversarial patches are robust to camera views, we show more images taken from different camera views in Figure 16. For the two cases: computer mouse (untargeted) and back bag (targeted to Yorkshire Terrier), we randomly sample ten other camera views. The results show that the adversarial patches are robust both for targeted settings and untargeted settings. For targteted settings, our adversarial patch can fool the network to predict back bag as terriers, and for untargeted settings, the adversarial patch misleads the network to predict computer mouse as artichoke.
@@
We presented the results when $\epsilon=16/255$ and found that: though designed for improved stealthiness and controllability, it is surprising that Diff-PGD transfers better than the original PGD to five different networks, here we show more results on different $\ell_{\infty}$ bound. We take the other two settings: $\epsilon=8/255$ and $\epsilon=32/255$ to better investigate our methods, for $\epsilon=8/255$ we still use $\eta=2/255$ while for $\epsilon=32/255$, we use $\eta=4/255$ for better attacks, both of them use $n=10$. For Diff-PGD, we use DDIM50 with $K_{s}=2$ for $\epsilon=8/255$ case and DDIM30 with $K_{s}=2$ for $\epsilon=32/255$ case.
@@
Figure 17: More Results on Transferability Test: in the main paper we show results of transferability of $\epsilon=16/255$ settings; here we test transferability of Diff-PGD under different $\ell_{\infty}$ bounds, (Left): $\epsilon=32/255,\eta=4/255,n=10$, for Diff-PGD we use DDIM50 with $K_{s}=2$, (Right): $\epsilon=8/255,\eta=2/255,n=10$ and for Diff-PGD we use DDIM30 with $K_{s}=2$. From the plots we can see that Diff-PGD has better transferability power than PGD.
Appendix H Human Evaluation
To better evaluate the stealthiness of Diff-PGD compared with the original PGD attack, we conduct a survey among humans with the assistance of Google Form. The user interface of the survey is shown in Figure 21, where the participants are allowed to choose multiple candidate images that they think are realistic. We randomly choose ten images from the ImageNet validation set and the choices for each question include one clean image, one adv-sample generated by PGD, and one adv-sample generated by Diff-PGD. For Diff-PGD we use DDIM50 with $K_{s}=3$, and for both Diff-PGD and PGD we use $\epsilon=16/255,\eta=2/255,n=10$.
@@ Appendix H Human Evaluation
Appendix I Potential Social Impacts
@@
Figure 19: More Results of Diff-PGD with $\mathbf{\epsilon=16/255}$: five columns of each image block follows $x$, $x_{PGD}$, $\delta_{PGD}$, $x^{n}_{0}$ and $\delta^{n}_{0}$ as is defined in the main paper.